OpenStudy (anonymous):

please, help any one , solve it 3^2x+3^(x+2) = 10/3

OpenStudy (anonymous):

\[3^{2x}+3^{x+2}=\frac{10}{3}\]?

OpenStudy (anonymous):

gimmick is to write \[3^{x+2}=3^2\times 3^x=9\times 3^x\] giving you a quadratic equation in \(3^x\) i.e. solve \[u^2+9u=\frac{10}{3}\] for \(u\) then replace \(u\) by \(3^x\)

OpenStudy (cggurumanjunath):